A family has $146,496, 30-year mortgage at 5.4% compounded monthly. Find the monthly payment. Also find the unpaid balance after the following periods of time. (A) 10 years (B) 20 years (C) 25 years.
Answer #1

Monthly payment=146496*(5.4%/12)/(1-1/(1+5.4%/12)^(12*30))=$822.62


Unpaid balance after 10 years=146496*(1+5.4%/12)^(12*10)-822.62/(5.4%/12)*((1+5.4%/12)^(12*10)-1)=$120,574.31


Unpaid balance after 20 years=146496*(1+5.4%/12)^(12*20)-822.62/(5.4%/12)*((1+5.4%/12)^(12*20)-1)=$76,146.36


Unpaid balance after 25 years=146496*(1+5.4%/12)^(12*25)-822.62/(5.4%/12)*((1+5.4%/12)^(12*25)-1)=$43,170.74